OpenStudy (anonymous):

Solve the system by the method of substitution. 2x-2y-z=-14 z=-8 2x+3y+3z=-6 A. (3, 8, –8) B. (–3, –8, –8) C. (–3, 8, 8) D. (–3, 8, –8)

OpenStudy (anonymous):

first substitute z=-8 in the first and second equation and you will have 2 equations in 2 unknowns

OpenStudy (anonymous):

You will get \[ 2 x - 2y =-22\\ 2 x+ 3y =18 \]

OpenStudy (anonymous):

the first equation gives \[ x-y =-11\\ x= y -11 \] Substitute x by its value to get\[ 2(y-11) + 3y =18\\ 2y -22 + 3y =18\\ 5y =40\\ y=8\\ x= y-11=8-11=-3 \] So it is D